The following data is available for Sheffield Repair Shop for 2019:

Repair technicians’ wages $290000
Fringe benefits 70000
Overhead 50000
Total

$410000


The desired profit margin is $34 per labor hour. The material loading charge is 40% of invoice cost. It is estimated that 5000 labor hours will be worked in 2019.

In March 2019, Sheffield repairs a bicycle that takes two hours to repair and uses parts of $180. The bill for this repair would be
Answer #1

Cost per labor hour = $ 410,000 /5000 = $ 82

For repair bill,

Cost of parts = $ 180

Material handling charge = 40% of invoice = 40% of 180 = $ 72

Total material cost for repair = 180 + 72 = $ 252

It takes two hours to repair

Cost per labor hour = $ 82

Desired margin = $ 34

labor cost per hour = 82+34 = $ 116

Labor cost for repair = 116*2 hours =$232

Total repair bill = material cost + labor cost = 252 + 232 = $ 484

Kindly upvote